# Revisiting electron-capture decay for Galactic cosmic-ray data

### v1

@@ 3.4 Validity of currently used approximations
The top panel of Fig. 5 shows $n$ at which the convergence is reached: above a few hundreds of MeV/n (IS), the 2-level approximation is sufficient for all EC-unstable species, but for a few nuclei with $Z\gtrsim 40$ and $t_{1/2}\gtrsim 10$ kyr. Nevertheless, at 10 MeV/n, the convergence is reached for $n=13$ levels for intermediate-lived nuclei of charge Z=3040Z=30-40. For the LBM, similar conclusions are reached (same number of levels to reach convergence), though with slightly tighter contours along the $t_{1/2}$ axis.
@@
The bottom panel shows the relative difference between the converged solution (assimilated to be $N_{13\,{\rm lev}}$) and $N_{2\,\rm lev}$. The flux is further suppressed when extra levels are present, and the largest for regions of large $n$ (compare the top and bottom panels). The relative difference between $N_{3\,\rm lev}$ and $N_{2\,\rm lev}$ (not shown) indicates that a large fraction of the extra-suppression is already contained in the $n=3$ level calculation, whereas each new level has a decreasing impact. This is not surprising, because going from 1 to 2 attached electrons doubles the EC-decay rate (two instead of one electron are available in the $K$-shell), whereas supplementary electrons (in other shells) do not enhance the decay and come at the cost of multiplying the overall attachment time.
@@
Among the isotopes for which CR data exist (i.e., those with the names highlighted in the plot), the converged solution differs at most by $\lesssim 20\%$, and only for very low energy (bottom left panel). More importantly, in the range targeted by TOA current experiments, that is, above several hundreds of MeV/n in equivalent IS energies (in-between the middle and right panels), the difference in the decayed fraction does not even reach the percent level, validating the use of the 2-level approximation. Similar conclusions are reached for the LBM.
@@
Table 2: CR isotopic fraction (weighted mean and weighted error) extracted from the following TOA data in CRDB: $^{7}{\rm Be}$/Be (Hagen et al., 1977; Garcia-Munoz et al., 1977; Webber et al., 1977; Buffington et al., 1978; Webber and Kish, 1979; Wiedenbeck and Greiner, 1980; Garcia-Munoz et al., 1981; Connell, 1998, 2001b; Nozzoli and Cernetti, 2021), $^{37}{\rm Ar}$/Ar (Young et al., 1981; Webber et al., 1985; Ogliore et al., 2009), $^{41}{\rm Ca}$/Ca (Young et al., 1981; Webber et al., 1985; Leske and Wiedenbeck, 1993; Connell, 2001b; Ogliore et al., 2009); $^{44}{\rm Ti}$/Ti, $^{49}{\rm V}$/V and $^{51}{\rm Cr}$/Cr (Young et al., 1981; Webber, 1981; Leske, 1993; Connell, 2001b); $^{53}{\rm Mn}$/Mn (Young et al., 1981; Webber, 1981; Leske, 1993; DuVernois, 1997; Connell, 2001b), $^{55}{\rm Fe}$/Fe (Young et al., 1981; Wiedenbeck et al., 2001; Connell, 2001b), $^{56}{\rm Ni}$/Ni (Young et al., 1981), $^{57}{\rm Co}$/Co (Leske, 1993; Connell, 2001b), $^{59}{\rm Ni}$/Ni (Connell, 2001b); $^{67}{\rm Ga}$/Ga, $^{71}{\rm Ge}$/Ge and $^{73}{\rm As}$/As (Binns et al., 2022). The dispersion between the data point values is much larger than the weighted errors for $20\leq Z\leq 26$ isotopes, and $Z\geq 28$ isotopes have only one data point.
$Z$ Isotopic ratio Weighed mean (error)
4 $^{7}{\rm Be}$/Be 0.57 ($\pm$ 0.06)
18 $^{37}{\rm Ar}$/Ar 0.03 ($\pm$ 0.02)
20 $^{41}{\rm Ca}$/Ca 0.07 ($\pm$ 0.02)
22 $^{44}{\rm Ti}$/Ti 0.01 ($\pm$ 0.003)
23 $^{49}{\rm V}$/V 0.40 ($\pm$ 0.05)
24 $^{51}{\rm Cr}$/Cr 0.24 ($\pm$ 0.02)
25 $^{53}{\rm Mn}$/Mn 0.44 ($\pm$ 0.03)
26 $^{55}{\rm Fe}$/Fe 0.04 ($\pm$ 0.02)
27 $^{57}{\rm Co}$/Co 0.32 ($\pm$ 0.20)
28 $^{56}{\rm Ni}$/Ni 0.04 ($\pm$ 0.03)
28 $^{59}{\rm Ni}$/Ni 0.02 ($\pm$ 0.02)
31 $^{67}{\rm Ga}$/Ga 0.07 ($\pm$ 0.04)
32 $^{71}{\rm Ge}$/Ge 0.11 ($\pm$ 0.07)
33 $^{73}{\rm As}$/As 0.36 ($\pm$ 0.23)
@@ 4 Detectability of EC decay in current data
Table 3: Detectability of EC-decay (in the 1D-DM) with current IS CR data. The first three columns show the charge, name, and half-life of the EC-unstable species (for which IS elemental flux data exist). The next three columns show the element name, energy range, and best precision reached by current elemental IS fluxes from Voyager 1 (Cummings et al., 2016). We then show the expected decayed fraction at 20 MeV/n IS energy for the isotopic (next-to-last column) and the elemental fluxes (last column). The latter is calculated from the isotopic one, accounting for the isotopic fractions reported in Table 2. We stress that both 56Ni and 59Ni contribute to the Ni flux.
@@ 4.1 EC-decay in IS data
Table 4: Energy range and best achieved precision on existing CR TOA for isotopic measurements (first three columns), either for EC-unstable parent-to-daughter ratios (Scott, 2005), or for isotopic ratios (same references as in Table 2).
@@
$Z$ Name $t_{1/2}$ Qty $E_{\rm k/n}$ [GeV/n] Prec. Qty $E_{\rm k/n}$ [GeV/n] Prec.
4 $^{7}{\rm Be}$ 53.22 d $^{7}{\rm Be}$$^{7}{\rm Li} - -$ $^{7}{\rm Be}$/Be [0.08-1.42] 2%
18 $^{37}{\rm Ar}$ 35.011 d $^{37}{\rm Ar}$$^{37}{\rm Cl}$ [0.12-0.35] 9% $^{37}{\rm Ar}$/Ar [0.49-0.57] 6%
20 $^{41}{\rm Ca}$ 99.4 kyr $^{41}{\rm Ca}$$^{41}{\rm K}$ [0.12-0.37] 10% $^{41}{\rm Ca}$/Ca [0.58-0.65] 6%
22 $^{44}{\rm Ti}$ 59.1 yr $^{44}{\rm Ti}$$^{44}{\rm Ca}$ [0.13-0.39] 14% $^{44}{\rm Ti}$/Ti [0.32-0.69] 27%
23 $^{49}{\rm V}$ 330 d $^{49}{\rm V}$$^{49}{\rm Ti}$ [0.14-0.41] 8% $^{49}{\rm V}$/V [0.32-0.69] 19%
24 $^{51}{\rm Cr}$ 27.7015 d $^{51}{\rm Cr}$$^{51}{\rm V}$ [0.14-0.42] 7% $^{51}{\rm Cr}$/Cr [0.32-0.72] 10%
25 $^{53}{\rm Mn}$ 3.7 Myr $^{53}{\rm Mn}$$^{53}{\rm Cr}$ [0.14-0.42] 9% $^{53}{\rm Mn}$/Mn [0.32-0.74] 10%
26 $^{55}{\rm Fe}$ 2.7562 yr $^{55}{\rm Fe}$$^{55}{\rm Mn}$ [0.14-0.43] 11% $^{55}{\rm Fe}$/Fe [0.62-0.94] 10%
28 $^{56}{\rm Ni}$ 6.075 d $^{56}{\rm Ni}$$^{56}{\rm Fe} - -$ $^{56}{\rm Ni}$/Ni [0.65-0.99] 74%
27 $^{57}{\rm Co}$ 271.811 d $^{57}{\rm Co}$$^{57}{\rm Fe}$ [0.15-0.44] 16% $^{57}{\rm Co}$/Co [0.32-0.32] 30%
28 $^{59}{\rm Ni}$ 81.82 kyr $^{59}{\rm Ni}$$^{59}{\rm Co} - -$ $^{59}{\rm Ni}$/Ni [0.10-0.50] 70%
31 $^{67}{\rm Ga}$ 3.2617 d $^{67}{\rm Ga}$$^{67}{\rm Zn} - -$ $^{67}{\rm Ga}$/Ga [0.13-0.70] 58%
32 $^{71}{\rm Ge}$ 11.43 d $^{71}{\rm Ge}$$^{71}{\rm Ga} - -$ $^{71}{\rm Ge}$/Ge [0.13-0.70] 64%
33 $^{73}{\rm As}$ 80.30 d $^{73}{\rm As}$$^{73}{\rm Ge} - -$ $^{73}{\rm As}$/As [0.13-0.70] 65%
@@ 4.2 From IS to TOA decaying fractions
    We assume that the variability of any IS flux shape is encompassed by the variability of He to Ni IS fluxes, as parametrised in Shen et al. (2019, 2025). We denote the corresponding fluxes by $\psi^{\rm IS}_{\rm Chen}$.
@@
The above relation implies that, for the same energy $f^{\rm TOA}\!<\!f^{\rm IS}$. It also means that $f^{\rm TOA}$ is larger for periods of low Solar activity than for high Solar acticity ones).
@@ 4.3 EC-decay in TOA isotopic data
EC-decay detectability
We show in Fig. 6 the comparison between the decayed fraction (lines) and the best data precision (filled rectangles). As before, we say that EC-decay is detectable if the decaying fraction of an isotope is larger than its data relative precision. First, light EC-unstable species are very rare, with only $^{7}{\rm Be}$ for $Z<15$. Even though the isotopic ratio of the latter is measured with the greatest precision ($\sim 2\%$), its decay fraction is less than 1% (owing to its long attachment time). Second, for $Z>15$, we have different behaviours depending on the isotopes half-life. The decay fraction for intermediate-lived isotopes is strongly dependent on the decay time: 41Ca ($t_{1/2}=99.4$ kyr), 53Mn ($t_{1/2}=3.7$ Myr) and 59Ni ($t_{1/2}=81.82$ kyr) have a decay fraction at the percent level (not shown) above a few hundreds of MeV/n, while 44Ti ($t_{1/2}=59.1$ yr) reaches 10% (solid cyan line). For short-lived isotopes ($t_{1/2}\lesssim$ yr), the decaying fraction strongly depends on the atomic number, with the heaviest species (67Ga, 71Ge and 73As) showing the highest decayed fraction.
@@
All in all, the detectability of EC-decay is possible if the curves are below the corresponding boxes (of same colour). This only clearly happens for 51Cr and 55Fe (dashed and dotted red lines), and otherwise, we barely meet the condition for 57Co (thin cyan line), and 67Ga, 71Ge and 73As (dashed, dotted, and dash-dotted blue lines). So it seems that we are at the verge of being able to detect EC-decay in CR data. However, this must be mitigated by the fact that EC-decay is overestimated in our modelling, as we neglected energy losses. And this is without considering the impact of nuclear production (Clinton and Waddington, 1993; Waddington, 1996, 1998) and inelastic cross-section uncertainties (Maurin et al., 2025), possibly larger than the decayed fraction itself.
@@ 4.4 EC decay in TOA elemental data
Elemental fluxes have been measured by a huge number of experiments. But at low energy and for $Z\leq 30$, the most precise data have a few percent precision, and are obtained down to $\simeq 100$ MeV/n from ACE-CRIS (Lave et al., 2013) and AMS-02666To be precise, AMS-02 data are in rigidity unit, and converting them in kinetic energy per nucleon introduces additional systematics at the few percent level (Derome et al., 2019), because the isotopic content is not measured above a few GeV/n. with a minimum energy of $\simeq 500$ MeV/n (Aguilar et al., 2021a, b). A summary view of the relevant data are gathered in Table 5, for the elemental fluxes (4th to 6th columns), but also for ratios of parent/daughter elemental fluxes (last 3 columns) – the latter ratios contain the EC-unstable isotope in their numerator and its daughter in their denominator, maximising the EC-decay impact. Only ratios have been measured for $Z>30$, by TIGER (Rauch et al., 2009), Super-TIGER (Murphy et al., 2016), ACE-CRIS (Binns et al., 2022) and CALET (Adriani et al., 2025). These experiments provide data up to $Z\simeq 40$, but that cannot be exploited here: the isotopic content of EC-unstable species is necessary to obtain the elemental decaying fraction, and this content has only been measured [^7] In principle, we could use the predicted isotopic content instead, but we do not have it in our very simplified approach. for Z=3133Z=31-33 (Binns et al., 2022).
@@
As discussed in the previous section, the TOA decayed fraction of intermediate-lives isotopes 7Be, 41Ca, 53Mn, and 59Ni are at the percent level, so when diluted by their isotopic content (0.57, 0.07, 0.44 and 0.02, respectively, see Table 2), this fraction is sub-percent in the elemental TOA fluxes (Be, Ca, Mn, and Ni, respectively). Even the decayed fraction in Ti is sub-percent, owing to its small isotopic fraction: the decayed fraction of 44Ti is 10%, diluted by its 0.01 isotopic fraction. Only short-lived EC-unstable isotopes provide $\gtrsim 1\%$ decayed fractions. We show, in Fig. 7, their expected decayed fractions (lines), compared to the CR data precision (rectangular boxes). For $Z\leq 30$, only V, Cr, and Co (thin-dashed, dotted, and thick-dashed blue lines, respectively) have data precision at the level or better than the expected decayed fraction, Co being the most favourable case. Actually, forthcoming AMS-02 data on these species will start at $\gtrsim 500$ MeV/n with a $\gtrsim 6\%$ (uncertainties dominated by systematics on the acceptance), so ACE-CRIS data are more favourable to search for this decay. For $Z>30$, neither Ga and Ge (red lines), nor As (cyan line) meet this requirement, and by a factor of two or more.

### v2

@@ 3.4 Validity of currently used approximations
[…]
@@
The bottom panel shows the relative difference between the converged solution (assumed to be $N_{13\,{\rm lev}}$) and $N_{2\,\rm lev}$. The flux is further suppressed when extra levels are present, and the largest for regions of large $n$ (compare the top and bottom panels). The relative difference between $N_{3\,\rm lev}$ and $N_{2\,\rm lev}$ (not shown) indicates that a large fraction of the extra-suppression is already contained in the $n=3$ level calculation, whereas each new level has a decreasing impact. This is not surprising, because going from 1 to 2 attached electrons doubles the EC-decay rate (two instead of one electron are available in the $K$-shell), whereas supplementary electrons (in other shells) do not enhance the decay and come at the cost of multiplying the overall attachment time.
@@
Among the isotopes for which CR data exist (i.e., those with the names highlighted in the plot), the converged solution differs at most by $\lesssim 20\%$ (apart from 59Ni showing a difference $\sim 37\%$), and only for very low energy (bottom left panel). More importantly, in the range targeted by TOA current experiments, that is, above several hundreds of MeV/n in equivalent IS energies (in-between the middle and right panels), the difference in the decayed fraction does not even reach the percent level, validating the use of the 2-level approximation. Similar conclusions are reached for the LBM.
@@
Table 2: CR isotopic fraction extracted from the following TOA data in CRDB: $^{7}{\rm Be}$/Be (Hagen et al., 1977; Garcia-Munoz et al., 1977; Webber et al., 1977; Buffington et al., 1978; Webber and Kish, 1979; Wiedenbeck and Greiner, 1980; Garcia-Munoz et al., 1981; Connell, 1998, 2001b; Nozzoli and Cernetti, 2021), $^{37}{\rm Ar}$/Ar (Young et al., 1981; Webber et al., 1985; Ogliore et al., 2009), $^{41}{\rm Ca}$/Ca (Young et al., 1981; Webber et al., 1985; Leske and Wiedenbeck, 1993; Connell, 2001b; Ogliore et al., 2009); $^{44}{\rm Ti}$/Ti, $^{49}{\rm V}$/V and $^{51}{\rm Cr}$/Cr (Young et al., 1981; Webber, 1981; Leske, 1993; Connell, 2001b); $^{53}{\rm Mn}$/Mn (Young et al., 1981; Webber, 1981; Leske, 1993; DuVernois, 1997; Connell, 2001b), $^{55}{\rm Fe}$/Fe (Young et al., 1981; Wiedenbeck et al., 2001; Connell, 2001b), $^{56}{\rm Ni}$/Ni (Young et al., 1981), $^{57}{\rm Co}$/Co (Leske, 1993; Connell, 2001b), $^{59}{\rm Ni}$/Ni (Connell, 2001b); $^{67}{\rm Ga}$/Ga, $^{71}{\rm Ge}$/Ge and $^{73}{\rm As}$/As (Binns et al., 2022). The last column shows the variance-weighted mean, $\langle x\rangle=(\sum_{i}w_{i}x_{i})/(\sum_{i}w_{i}$) with $w_{i}=1/\sigma_{i}^{2}$ – which corresponds to the maximum likelihood estimator for independent and normally distributed values – and its standard error ($\sigma_{\langle x\rangle}=\sqrt{1/(\sum_{i}w_{i})}$, calculated over all energies of the CR data available. We stress that the
dispersion between the data points is much larger than the weighted errors for $20\leq Z\leq 26$ isotopes, and that $Z\geq 28$ isotopes have only one data point. $Z$ Isotopic ratio Weighed mean (error) 4 $^{7}{\rm Be}$/Be 0.57 ($\pm$ 0.06) 18 $^{37}{\rm Ar}$/Ar 0.03 ($\pm$ 0.02) 20 $^{41}{\rm Ca}$/Ca 0.07 ($\pm$ 0.02) 22 $^{44}{\rm Ti}$/Ti 0.01 ($\pm$ 0.003) 23 $^{49}{\rm V}$/V 0.40 ($\pm$ 0.05) 24 $^{51}{\rm Cr}$/Cr 0.24 ($\pm$ 0.02) 25 $^{53}{\rm Mn}$/Mn 0.44 ($\pm$ 0.03) 26 $^{55}{\rm Fe}$/Fe 0.04 ($\pm$ 0.02) 27 $^{57}{\rm Co}$/Co 0.32 ($\pm$ 0.20) 28 $^{56}{\rm Ni}$/Ni 0.04 ($\pm$ 0.03) 28 $^{59}{\rm Ni}$/Ni 0.02 ($\pm$ 0.02) 31 $^{67}{\rm Ga}$/Ga 0.07 ($\pm$ 0.04) 32 $^{71}{\rm Ge}$/Ge 0.11 ($\pm$ 0.07) 33 $^{73}{\rm As}$/As 0.36 ($\pm$ 0.23)
@@ 4 Detectability of EC decay in current data
Table 3: Detectability of EC-decay (in the 1D-DM) with current IS CR data. The first three columns show the charge, name, and half-life of the EC-unstable species (for which IS elemental flux data exist). The next three columns show the element name, energy range, and best precision reached by current elemental IS fluxes from Voyager 1 (Cummings et al., 2016). We then show the expected decayed fraction computed from the 1D-DM model at 20 MeV/n IS energy for the isotopic (next-to-last column) and the elemental fluxes (last column). The latter is calculated from the isotopic one, accounting for the isotopic fractions reported in Table 2. We stress that both 56Ni and 59Ni contribute to the Ni flux.
@@ 4.1 EC-decay in IS data
Table 4: Energy range, variance-weighted mean, and best achieved precision on existing CR TOA data ratios involving EC-unstable species: parent-to-daughter ratios (Scott, 2005) and isotopic abundances (same references and weighted-mean already reported in Table 2).
@@
$Z$ Name $t_{1/2}$ Qty $E_{\rm k/n}$ [GeV/n] Mean Prec. Qty $E_{\rm k/n}$ [GeV/n] Prec.
4 $^{7}{\rm Be}$ 53.22 d $^{7}{\rm Be}$$^{7}{\rm Li} - - -$ $^{7}{\rm Be}$/Be [0.08-1.42] 2%
18 $^{37}{\rm Ar}$ 35.011 d $^{37}{\rm Ar}$$^{37}{\rm Cl}$ [0.12-0.35] 1.40 9% $^{37}{\rm Ar}$/Ar [0.49-0.57] 6%
20 $^{41}{\rm Ca}$ 99.4 kyr $^{41}{\rm Ca}$$^{41}{\rm K}$ [0.12-0.37] 0.855 10% $^{41}{\rm Ca}$/Ca [0.58-0.65] 6%
22 $^{44}{\rm Ti}$ 59.1 yr $^{44}{\rm Ti}$$^{44}{\rm Ca}$ [0.13-0.39] 0.048 14% $^{44}{\rm Ti}$/Ti [0.32-0.69] 27%
23 $^{49}{\rm V}$ 330 d $^{49}{\rm V}$$^{49}{\rm Ti}$ [0.14-0.41] 3.10 8% $^{49}{\rm V}$/V [0.32-0.69] 19%
24 $^{51}{\rm Cr}$ 27.7015 d $^{51}{\rm Cr}$$^{51}{\rm V}$ [0.14-0.42] 2.52 7% $^{51}{\rm Cr}$/Cr [0.32-0.72] 10%
25 $^{53}{\rm Mn}$ 3.7 Myr $^{53}{\rm Mn}$$^{53}{\rm Cr}$ [0.14-0.42] 3.60 9% $^{53}{\rm Mn}$/Mn [0.32-0.74] 10%
26 $^{55}{\rm Fe}$ 2.7562 yr $^{55}{\rm Fe}$$^{55}{\rm Mn}$ [0.14-0.43] 1.52 11% $^{55}{\rm Fe}$/Fe [0.62-0.94] 10%
28 $^{56}{\rm Ni}$ 6.075 d $^{56}{\rm Ni}$$^{56}{\rm Fe} - - -$ $^{56}{\rm Ni}$/Ni [0.65-0.99] 74%
27 $^{57}{\rm Co}$ 271.811 d $^{57}{\rm Co}$$^{57}{\rm Fe}$ [0.15-0.44] 0.062 16% $^{57}{\rm Co}$/Co [0.32-0.32] 30%
28 $^{59}{\rm Ni}$ 81.82 kyr $^{59}{\rm Ni}$$^{59}{\rm Co} - - -$ $^{59}{\rm Ni}$/Ni [0.10-0.50] 70%
31 $^{67}{\rm Ga}$ 3.2617 d $^{67}{\rm Ga}$$^{67}{\rm Zn} - - -$ $^{67}{\rm Ga}$/Ga [0.13-0.70] 58%
32 $^{71}{\rm Ge}$ 11.43 d $^{71}{\rm Ge}$$^{71}{\rm Ga} - - -$ $^{71}{\rm Ge}$/Ge [0.13-0.70] 64%
33 $^{73}{\rm As}$ 80.30 d $^{73}{\rm As}$$^{73}{\rm Ge} - - -$ $^{73}{\rm As}$/As [0.13-0.70] 65%
@@ 4.2 From IS to TOA decaying fractions
    We assume that the range of possible IS flux shapes is encompassed by the shapes of He to Ni IS fluxes, as parametrised in Shen et al. (2019, 2025). We denote the corresponding fluxes by $\psi^{\rm IS}_{\rm Chen}$.
@@
The above relation implies that, for the same energy $f^{\rm TOA}\!<\!f^{\rm IS}$. It also means that $f^{\rm TOA}$ is larger for periods of low Solar activity than for high Solar activity ones.
@@ 4.3 EC-decay in TOA isotopic data
EC-decay detectability in parent nuclei
We show in Fig. 6 the comparison between the decayed fraction computed in the 1D-DM model (lines) and the best data precision (filled rectangles). As before, we say that EC-decay is detectable if the decaying fraction of an isotope is larger than its data relative precision. First, light EC-unstable species are very rare, with only $^{7}{\rm Be}$ for $Z<15$. Even though the isotopic ratio of the latter is measured with the greatest precision ($\sim 2\%$), its decay fraction is less than 1% (owing to its long attachment time). Second, for $Z>15$, we have different behaviours depending on the isotopes half-life. The decay fraction for intermediate-lived isotopes is strongly dependent on the decay time: 41Ca ($t_{1/2}=99.4$ kyr), 53Mn ($t_{1/2}=3.7$ Myr) and 59Ni ($t_{1/2}=81.82$ kyr) have a decay fraction at the percent level (not shown) above a few hundreds of MeV/n, while 44Ti ($t_{1/2}=59.1$ yr) reaches 10% (solid cyan line). For short-lived isotopes ($t_{1/2}\lesssim$ yr), the decaying fraction strongly depends on the atomic number, with the heaviest species (67Ga, 71Ge and 73As) showing the highest decayed fraction.
@@
All in all, the detectability of EC-decay is possible if the curves are below the corresponding boxes (of same colour). This only clearly happens for 51Cr and 55Fe (dashed and dotted red lines), and otherwise, we barely meet the condition for 57Co (thin cyan line), and 67Ga, 71Ge and 73As (dashed, dotted, and dash-dotted blue lines). So it seems that we are at the verge of being able to detect EC-decay in TOA CR data. However, this must be mitigated by the fact that EC-decay is overestimated in our modelling, as we neglected energy losses. And this is without considering the impact of nuclear production (Clinton and Waddington, 1993; Waddington, 1996, 1998) and inelastic cross-section uncertainties (Maurin et al., 2025), possibly larger than the decayed fraction itself.
EC-decay detectability in daughter nuclei
EC decay could also be sought in the daughter of the EC-unstable species. The weighted-average mean of the parent-to-daughter ratio (denoted $r$ in this paragraph), reported in Table 4 (6th column), gives an indication whether this would be more favourable than looking at the parent. Indeed, the daughter EC-fed fraction, $f_{d}$, is simply linked to the EC-decayed fraction of the parent, $f_{p}$, by $f_{d}=r\times f_{p}$. For ratios with $r<1$ (e.g., 41Ca/41K), EC decay will be even harder to see in the daughter than in the parent EC-unstable species. This is the reverse for ratios with $r>1$, and five ratios meet this criterion in Table 4. We provide in Table 6 their $f_{p}$ and $f_{d}$ values: the already spotted 49V/49Ti and 51Cr/51V ratios become even more favourable than before, thanks to the increased EC-decay impact in the daughter (in the denominator). In addition, 55Fe/55Mn now passes the detectability criterion, with $f_{d}$ larger than the CR data precision (rel. err. column in the Table). These trends would need to be confirmed by a complete analysis involving the full nuclear network. Indeed, Ti, V, and Mn isotopes are of secondary origin, so that they are maximally sensitive to the cross-section uncertainties mentioned earlier.
Table 6: Impact of EC-decay in parent-to-daughter ratios, for cases in which the detectability is larger in the daughter than in the parent. The first three columns (directly reported from Table 4) correspond to the CR TOA data ratio, weighted-average (denoted $r$), and best data precision. The next-to-last column reports the decayed fraction $f_{p}$ in the parent, as read off Fig. 6, at the most favourable energy $\sim 150$ MeV/n where data exist. The last column reports the decayed fraction fed to the daughter, $f_{d}=r\times f_{p}$, at this same most favourable energy
. Ratio $r$ rel.err. $f_{p}$ $f_{d}$ 37Ar/37Cl 1.40 9% 6% 8.4% 49V/49Ti 3.10 8% 12% 37.2% 51Cr/51V 2.52 7% 16% 40.3% 53Mn/53Cr 3.60 9% $1\% <$$3.6\%$ 55Fe/55Mn 1.52 11% 20% 30.4%
@@ 4.4 EC decay in TOA elemental data
Elemental fluxes have been measured by a huge number of experiments. But at low energy and for $Z\leq 30$, the most precise data have a few percent precision, and are obtained down to $\simeq 100$ MeV/n from ACE-CRIS (Lave et al., 2013) and AMS-02666To be precise, AMS-02 data are in rigidity unit, and converting them in kinetic energy per nucleon introduces additional systematics at the few percent level (Derome et al., 2019), because the isotopic content is not measured above a few GeV/n. with a minimum energy of $\simeq 500$ MeV/n (Aguilar et al., 2021a, b). A summary view of the relevant data is gathered in Table 5, for the elemental fluxes (4th to 6th columns), but also for ratios of parent/daughter elemental fluxes (last 3 columns) – the latter ratios contain the EC-unstable isotope in their numerator and its daughter in their denominator, maximising the EC-decay impact. Only ratios have been measured for $Z>30$, by TIGER (Rauch et al., 2009), Super-TIGER (Murphy et al., 2016), ACE-CRIS (Binns et al., 2022) and CALET (Adriani et al., 2025). These experiments provide data up to $Z\simeq 40$, but that cannot be exploited here: the isotopic content of EC-unstable species is necessary to obtain the elemental decaying fraction, and this content has only been measured [^7] In principle, we could use the predicted isotopic content instead, but we do not have it in our very simplified approach. for Z=3133Z=31-33 (Binns et al., 2022).
@@
As discussed in the previous section, the TOA decayed fraction of intermediate-lived isotopes 7Be, 41Ca, 53Mn, and 59Ni are at the percent level, so when diluted by their isotopic content (0.57, 0.07, 0.44 and 0.02, respectively, see Table 2), this fraction is sub-percent in the elemental TOA fluxes (Be, Ca, Mn, and Ni, respectively). Even the decayed fraction in Ti is sub-percent, owing to its small isotopic fraction: the decayed fraction of 44Ti is 10%, diluted by its 0.01 isotopic fraction. Only short-lived EC-unstable isotopes provide $\gtrsim 1\%$ decayed fractions. We show, in Fig. 7, their expected decayed fractions (lines), compared to the CR data precision (rectangular boxes). For $Z\leq 30$, only V, Cr, and Co (thin-dashed, dotted, and thick-dashed blue lines, respectively) have data precision at the level or better than the expected decayed fraction, Co being the most favourable case. Actually, forthcoming AMS-02 data on these species will start at $\gtrsim 500$ MeV/n with a $\gtrsim 6\%$ (uncertainties dominated by systematics on the acceptance), so ACE-CRIS data are more favourable to search for this decay. For $Z>30$, neither Ga and Ge (red lines), nor As (cyan line) meet this requirement, and by a factor of two or more.